OpenStudy (anonymous):

how do i figure out what percent of 160 is 24?

OpenStudy (anonymous):

Never heard a question like that before. But what i guess it means can be represented mathematically like this: (160x)/100 = 24 So 160x = 2400 x = 2400/160 = 15. So 15% of 160 is 24.

OpenStudy (anonymous):

ok, so it's like going backwards. thx!

OpenStudy (anonymous):

Yeah, thats exactly what it is